what mass of bromine will be produced if a potassium bromide solution reacts with 60.0mL of a 1.30mol/L solution of acidified NaCIO4


1
Expert's answer
2021-02-01T03:52:30-0500

Solution:

The balanced chemical equation:

4H2SO4 + NaClO4 + 8KBr → NaCl + 4H2O + 4Br2 + 4K2SO4

According to the equation above: n(NaClO4) = n(Br2)/4

Moles of NaClO4 = Molarity of NaClO4 × Solution volume

Moles of NaClO4 = 1.30 mol/L × 0.06 L = 0.078 mol


n(Br2) = 4 × n(NaClO4) = 4 × 0.078 mol = 0.312 mol


Moles of Br2 = Mass of Br2 / Molar mass of Br2

Mass of Br2 = Moles of Br2 × Molar mass of Br2

The molar mass of Br2 is 159.808 g mol-1.

Hence,

Mass of Br2 = 0.312 mol × 159.808 g mol-1 = 49.86 g

Mass of Br2 = 49.86 g


Answer: 49.86 grams of bromine (Br2) will be produced.